sam wants to invest part of his $24000 lottery winnings in an international stock fund that yields 4% annual interest and a domestic money market fund that earns 6.5% annual interest. find how much is invested in each account to earn $1360 interest at the end of one year. (use the formula, i=prt for simple, where i is the interest earned, p is the principal, r is the yearly interest rate, and t is the time in years.)
Answer by addingup(3677)

0.04x + 0.065(24000-x) = 1360
0.04x + 1560-0.065x = 1360
-0.025x = -200 Divide both sides by 0.025 and remember -/- = +
x = 8000 is the amount invested at 4%, and the amount at 6.5%:
24000-8000 = 16000
:
John